This is a 5th grade math problem that I heard from a friend and I am curious to learn the answer.
A student has a test with 2 parts. Part A questions are worth 2 points and Part B are worth 5 points each. The student gets all the questions right in part A but misses 1 question in part B. The student scores an 85% on the overall test. How many questions are on the test? Help I tried to do:
2x__ +5x___ (-5) = .85 but I can't make it come out to 85%
Thanks for your help

Answer by stanbon(75887)

A student has a test with 2 parts.
Part A questions are worth 2 points and Part B are worth 5 points each.
The student gets all the questions right in part A but misses 1 question in part B.
-------------------------
The student scores an 85% on the overall test. How many questions are on the test? Help I tried to do:
2x__ +5x___ (-5) = .85 but I can't make it come out to 85%
===========================================================
1st part:
x questions
2x points
----
2nd part:
y questions
5(y-1)
---
Average = (total points)/(# of questions)
[(2x)+5y-5]/(x+y) = 0.85
2x + 5y - 5 = 0.85x + 0.85y
1.15x + 4.15y = 5
------
Comment: Need to know more about x and y to solve.
=====================
Cheers,
Stan H.